2x+(k-1)y=6
3x+(2k+1)y=9

Answer by vidhyak(98)

2x + (k-1)y =6 ........eq1
3x + (2k+1)y =9 ........eq2

Multiply eq1 by 3 and eq2 by 2 and subtract
6x + 3(k-1)y = 18
-6x - 2(2k+1)y = -18

3(k-1)y - 2(2k+1)y = 0
y(3(k-1) - 2(2k+1)) = 0
3(k-1) - 2(2k+1) = 0
3k - 3 - 4k - 2 = 0
-k -5 = 0
-k = 5
k = -5